### v1

# Advancing Multi-Organ Disease Care: A Hierarchical Multi-Agent Reinforcement Learning Framework

Abstract
Multi-organ diseases present significant challenges due to their simultaneous impact on multiple organ systems, necessitating complex and adaptive treatment strategies. Despite recent advancements in AI-powered healthcare decision support systems, existing solutions are limited to individual organ systems. They often ignore the intricate dependencies between organ system and thereby fails to provide holistic treatment recommendations that are useful in practice. We propose a novel hierarchical multi-agent reinforcement learning (HMARL) framework to address these challenges. This framework uses dedicated agents for each organ system, and model dynamic through explicit inter-agent communication channels, enabling coordinated treatment strategies across organs. Furthermore, we introduce a dual-layer state representation technique to contextualize patient conditions at various hierarchical levels, enhancing the treatment accuracy and relevance. Through extensive qualitative and quantitative evaluations in managing sepsis—a complex multi-organ disease—our approach demonstrates its ability to learn effective treatment policies that significantly improve patient survival rates. This framework marks a substantial advancement in clinical decision support systems, pioneering a comprehensive approach for multi-organ treatment recommendations.
Introduction
Multi-organ diseases are characterized by the sequential or simultaneous impairment of multiple organ systems (Asim, Amin, and El-Menyar 2020). They present significant challenges in clinical management due to their complexity, the difficulty of balancing therapeutic trade-offs, and the potential for life-threatening outcomes in critically ill patients. Treating these diseases requires a holistic approach that accounts for the interdependencies between different organ systems (Tian et al. 2023). Existing guideline-based approaches treat organ systems in isolation and rely on one-size-fits-all recommendations (Whelehan, Conlon, and Ridgway 2020), which fail to address the complexities of multi-organ diseases. One example of such a disease is COVID-19, which, while primarily affecting the respiratory system, can also lead to dysfunction in the immune, nervous, and gastrointestinal systems (Thakur et al. 2021; Bhadoria and Rathore 2021). Another example is sepsis, a serious condition triggered by the body’s dysregulated response to infection. Sepsis can lead to widespread inflammation, coagulation abnormalities, and metabolic disruptions, cascading into multi-organ dysfunction, which requires comprehensive and adaptive treatment strategies (Greco et al. 2017).
Recent advances in artificial intelligence, particularly in reinforcement learning (RL), have shown promise in optimizing clinical decision-making for complex diseases. RL’s capacity to learn adaptive policies from high-dimensional, complex data makes it a powerful tool for treatment recommendations. Notably, Komorowski et al., pioneered the use of deep RL for sepsis treatment, developing a model that learns optimal treatment policies from electronic health records (EHRs) of intensive care unit (ICU) patients (Komorowski et al. 2018). Subsequent work in this area has explored new model-free approaches, such as those based on Dueling Double Deep Q-Networks (D3QN) (Raghu et al. 2017; Wu et al. 2023), as well as model-based approaches (Raghu, Komorowski, and Singh 2018). Beyond sepsis, RL has also been applied to chronic disease management, such as in the work by Zheng et al. (Zheng et al. 2021), which proposed an RL-based method for personalized diabetes and multimorbidity management by modeling health outcomes like glycemia, blood pressure, and cardiovascular disease risk. Liu et al., trained an RL model to recommend the dosage of oral antidiabetic drugs and insulin (Liu et al. 2020b). Additionally, Wang et al., proposed a model-based RL framework consisting of a patient model paired with a policy model to optimize insulin regimens through the analysis of glycemic response. (Wang et al. 2023).
Despite the multi-organ nature of many major diseases, existing RL approaches have predominantly focused on recommending treatments targeting a single organ system. For example, current sepsis-targeted solutions primarily address only the cardiovascular system, neglecting other relevant organ systems (Liu et al. 2020a). This is a significant limitation, as treatments for one organ system can significantly influence the efficacy or safety of treatments for another. For instance, recommending vasopressors (VAs) to stabilize blood pressure may not be feasible in a patient with concurrent renal dysfunction, as doing so could increase renal impairment (Yagi et al. 2021). While there have been AI-based solutions developed with explicit consideration of multiple organ systems, these solutions focus on the task of diagnosis, instead of the more complex treatment recommendation task (Khan et al. 2023; Kaur and Kaur 2024).
Multi-organ disease treatment recommendation tasks introduce complexities beyond that which traditional RL and non-RL-based recommendation system can effectively manage. For example, as the number of considered organ systems or treatments increase, the number of total possible combinations of actions can increase exponentially, quickly making action spaces untenable for standard single-agent RL algorithms to navigate. Additionally, patients’ measured physiological variables will relate to each organ system’s unique physiology in different ways, adding an additional layer of complexity that a multi-organ solution must account for. Consequently, there is a strong need to develop a robust and holistic treatment recommendation system tailored for multi-organ disease management.
To this end, we propose a hierarchical multi-agent RL (HMARL) system. In this framework, the complex task of multi-organ treatment recommendation is divided among a hierarchy of specialized sub-agents. Each agent operates within its own localized state and action spaces, simplifying decision-making and allowing agents to focus exclusively on relevant subspaces. Carefully designed inter- and intra- agent communication mechanisms enable collaboration when necessary, alleviating the burden on individual agents and leading to more efficient training and faster convergence to optimal policies. Moreover, understanding patient states and their dynamics within localized contexts is essential for accurate treatment recommendations. For example, when treating the cardiovascular system, more focus should be on factors such as ejection fraction and cardiac enzyme levels, whereas renal treatments require attention to factors like glomerular filtration rate and electrolyte balances (Deferrari, Cipriani, and La Porta 2021). To achieve this, we propose a multi-layer hierarchical representation technique that first captures broad health indicators at the root level and then refines them into organ-specific representations, which are utilized by the corresponding agents at these levels. Collectively, this HMARL system, combined with the multi-layer hierarchical representation technique, effectively manages the complexities of multi-organ treatment recommendation.
We summarize our main contributions as follows:
  • To the best of our knowledge, we propose the first-of-its-kind multi-organ treatment recommendation solution.
  • We introduce a compact HMARL framework that decomposes the complex task of multi-organ disease management into manageable sub-tasks, handled by specialized sub-agents operating within localized state and action spaces, both independently and collaboratively.
  • We develop a multi-layer hierarchical representation technique that learns both broad and specific patient representations, tailored to treatment context, aiding accurate decision-making at multiple levels.
  • We demonstrate the effectiveness of our approach through extensive experiments, showing its superiority over traditional RL models in handling multi-organ interdependencies and improving treatment outcomes.
It must be emphasized that our solution—trained and tested on retrospective public data—serves purely as a clinical decision support system. It is not intended to take on a decision-making role, but merely offers data-driven treatment recommendations to an expert-in-the-loop in the clinical setting. There are no ethical violations in the development or use of this solution.
Methodology
Hierarchical Decomposition
Figure 1: HMARL solution architecture.
We decompose the intricate multi-organ recommendation task into a clinically meaningful hierarchy of sub-tasks, recommending seven treatments across three organ systems ( Figure 1). This structured decomposition significantly reduces the overall complexity by confining each sub-task to its own subspace, only coordinating with other subtasks when necessary.
At the top level, the root agent ($M_{Rt}$) selects from five primary actions: No-treatment ($\mathcal{A}_{Rt}^{0}$), Neuro-only ($\mathcal{A}_{Rt}^{Neu}$), Cardio-only ($\mathcal{A}_{Rt}^{Car}$), Renal only ($\mathcal{A}_{Rt}^{Ren}$) or Organ mixture ($\mathcal{A}_{Rt}^{OMix}$). Including the frequently utilized $\mathcal{A}_{Rt}^{0}$ option at the root level substantially simplifies the process by reducing the overall action space and eliminating the need to consider this non-intervention option in subsequent tasks.
Organ-specific actions ($\mathcal{A}_{Rt}^{Neu}$, $\mathcal{A}_{Rt}^{Car}$, $\mathcal{A}_{Rt}^{Ren}$) invoke respective agents ($M_{Neu}$, $M_{Car}$, $M_{Ren}$), each operating within dedicated treatment subspaces. Subsequently, $M_{Neu}$ chooses from four possible actions: S1-only ($\mathcal{A}_{Neu}^{S1}$), S2-only ($\mathcal{A}_{Neu}^{S2}$), S3-only ($\mathcal{A}_{Neu}^{S3}$), or a mixture of them ($\mathcal{A}_{Neu}^{Mix}$), each invoking a specialized sub-agent ($M_{Neu}^{S1}$, $M_{Neu}^{S2}$, $M_{Neu}^{S3}$ or $M_{Neu}^{Mix}$), recommending individual or combined treatments. $M_{Car}$ follows a similar structure, selecting from three options: IV-only ($\mathcal{A}_{Car}^{IV}$), Vaso-only ($\mathcal{A}_{Car}^{VA}$), and IV and Vaso mixture ($\mathcal{A}_{Car}^{Mix}$), invoking dedicated sub-agents ($M_{Car}^{IV}$, $M_{Car}^{VA}$ or $M_{Car}^{Mix}$), recommending individual or treatment mixtures. The recommendations for renal agent $M_{Ren}$ is comparatively simpler. Since there is no mixing between the treatments (diuretics and dialysis), $M_{Ren}$ directly chooses from four diuretic levels or. When recommending dosage mixtures for a single organ, respective sub-agents collaborate to provide a coordinated recommendation. Each treatment-specific sub-agent communicates its recommendations to the dedicated mixture agent ($M_{Neu}^{S1}$, $M_{Neu}^{S2}$, $M_{Neu}^{S3}$ to $M_{Neu}^{Mix}$ and $M_{Car}^{IV}$, $M_{Car}^{VA}$ to $M_{Car}^{Mix}$), if they were to be invoked. This additional intra-organ communication allows the mixture agent to integrate insights from respective sub-agents.
When complex multi-organ treatments are recommended ($\mathcal{A}_{Rt}^{OMix}$), the $M_{OMix}$ agent is invoked. It consults with organ-specific agents ($M_{Neu}$, $M_{Car}$, $M_{Ren}$) through inter-organ communications to access their dosage recommendations (if they were to be invoked). These auxiliary inputs help $M_{OMix}$ comprehensively understand the possible treatments and their scopes, greatly reducing its decision-making burden. Subsequently, $M_{OMix}$ invokes a combination of sub-agents (at least two pairs of $M_{OMix}^{Neu}$, $M_{OMix}^{Car}$ and $M_{OMix}^{Ren}$ agents), where these sub-agents each communicate with the opposite agents from the single-organ subtasks. For example, $M_{OMix}^{Neu}$ considers the leaf level recommendations from $M_{Car}$ and $M_{Ren}$’s sub-agents as additional information. This system benefits the agents by allowing them to adjust communicated recommendations based on their use in combination with other treatments, rather than needing to devise them from scratch. Consequently, this facilitates the agents’ ability to efficiently determine the optimal treatment combination, streamlining the decision-making process for complex, multi-organ interventions. All agents are trained exclusively on samples from their specific action subspaces. For instance, the $M_{Car}^{IV}$ agent is trained on cases where IV was the only treatment administered, while $M_{Car}^{Mix}$ agent is trained on samples involving simultaneous use of both IV and Vaso treatments.
This approach mimics a collaborative clinical setting where a lead physician oversees specialists, each prescribing treatments in their expertise to ensure comprehensive and specialized care. By distributing recommendation tasks among specific and mixture agents, our structured hierarchy efficiently handles the complexities of multi-treatment recommendations within and across organ systems. The decomposition into precise subtasks is determined by factors such as the number of treatments and their combinations. The carefully designed communication channels enables collaboration and effectively employ a divide-and-conquer strategy. We believe that the proposed hierarchical decomposition approach is both flexible and scalable, thereby making it highly suitable for various medical scenarios involving multi-organ treatments, beyond this use case.
RL Components
RL States:
Accurate patient understanding is essential for effective multi-organ treatment recommendations, which require integrating complex interdependencies of physiological features and organ functions. To address this, we propose a hierarchical patient representation approach that acknowledges the varying significance of raw features at different analytical levels.
At the root level, Unified State Representations are learned to extract broad health indicators and their dynamics, providing a foundational understanding of patient status. This representation is used for broader decision-making processes at the root level, and sets the stage for subsequent, more granular recommendations. At the organ levels, they are refined to learn Targeted State Representations, tailored to unique physiological requirements and and interrelationships of specific organs. For instance, in cardiac treatments, the embeddings prioritize features like ejection fraction and cardiac enzyme levels, capturing essential cardiac health indicators. In contrast, renal treatments focus on features like glomerular filtration rate and electrolyte balances, crucial for assessing renal function. This approach ensures that organ-specific recommendations are precise and relevant.
This dual-layer hierarchical representation strategy balances broad applicability with detailed specificity, enhancing decision-making capabilities by considering each organ’s condition within the overall health context. The hierarchical structure comprises the following levels:
Unified State Representations: Inspired by the representation learning proposed by Perera et al., (Perera, Liu, and Feng 2023) we learn the unified representations as follows. Each patient at time $t$ is represented by their raw $d$-dimensional feature, $x_{t}=\{x_{t,1},x_{t,2},\ldots,x_{t,d}\}\in\mathbb{R}$ (for details on $x_{t}$, see Appendix Section Feature Processing ). At the root agent level, these features are transformed using dense latent embeddings $E^{Rt}=\{e_{1}^{Rt},e_{2}^{Rt},\cdots,e_{d}^{Rt}\}\in\mathbb{R}^{d\times k}$. Each $k$-dimensional latent embedding vector $e_{i}^{Rt}\in\mathbb{R}^{k}$ transforms its corresponding raw feature into a more informative dense latent representation. The resultant patient-specific embeddings are represented as $F^{Rt}_{t}=\{f_{t,1}^{Rt},f_{t,2}^{Rt},\cdots,f_{t,d}^{Rt}\}\in\mathbb{R}^{d\times k}$, where $f_{t,i}^{Rt}=(x_{t,i}\cdot e_{i}^{Rt})$. These latent embeddings $e_{i}^{Rt}$ are generic at the root agent level, providing a holistic understanding of the patient by transforming each feature into a homogeneous latent space. The vector $f_{t,i}^{Rt}$ denotes the transformed representation specific to the patient at time $t$.
To capture the complex interdependencies among these features, a higher-order interaction layer is introduced. This layer computes the element-wise product between all pairs of embeddings in $F_{t}^{Rt}$ resulting an interaction matrix $G_{t}^{Rt}\in\mathbb{R}^{k\times d(d+1)/2}$. Final output of the layer consists of both first- and second- order interactions, denoted by $H_{t}^{Rt}=(F_{t}^{Rt}\mid G_{t}^{Rt})$, which is then aggregated via sum pooling to generate an observation vector as $o_{t}^{Rt}=\sum_{l=1}^{d(d+3)/2}H_{t,l}^{Rt}\in\mathbb{R}^{k}$.
Moreover, given the importance of patient trajectory information for understanding the patient’s current context, a temporal contextual state vector $c_{t}^{Rt}$. This vector captures the recent history of the patient’s states by applying an exponential decay to the previous observation vectors, resulting in $c_{t}^{Rt}=\sum_{i=t-3}^{t-1}e^{-(t-i)}o_{i}^{Rt}\in\mathbb{R}^{k}$.
The final core-state vector at the root level at time $t$ is constructed by concatenating the current observation vector $o_{Rt}$ with the temporal contextual state vector $c_{Rt}$, resulting in $s_{t}^{Rt}=(o_{t}^{Rt}\mid c_{t}^{Rt})\in\mathbb{R}^{2k}$. This vector is learned end-to-end during the training of the root agent, enabling the model to capture both immediate and historical physiological measurements effectively. The reduced dimensionality to $2k$ in the dense space also simplifies the complexity of subsequent decision-making processes.
Targeted State Representations: The targeted (organ-level) state representations are developed by fine-tuning the generic latent embeddings $E^{Rt}$ learned at the root level to reflect the unique characteristics and critical features relevant to each organ. Accordingly, using a similar approach, we transform the raw features $x_{t}$ using specialized latent embeddings $E^{Neu}$, $E^{Car}$ and $E^{Ren}$, to obtain organ-specific state representations $s_{t}^{Neu}$, $s_{t}^{Car}$ and $s_{t}^{Ren}$. They are used as input states for training the corresponding sub-agents ($M_{Neu}$, $M_{Car}$ and $M_{Ren}$). Moreover, a concatenated targeted state representation $s_{t}^{OMix}=\left[s_{t}^{Neu}|s_{t}^{Car}|s_{t}^{Ren}\right]$ is used for training all organ mixture agents ($M_{OMix}$, $M^{Neu}_{OMix}$, $M^{Car}_{OMix}$ and $M^{Ren}_{OMix}$), ensuring they are trained on a comprehensive representation of all targeted organs.
RL Actions:
In addition to the agent options discussed above for higher-level agents, this section provides detailed information on the actions of leaf-level agents. According to our hierarchical task decomposition approach, each agent operates within a factored action space. It helps reduce the learning burden on each agent and prevents the effects of potentially low-quality samples across subspaces. Based on clinical expertise, we selected seven treatments. Two treatments were considered for the Cardiovascular system: IV fluids (IV) and vasopressors (Vaso), three for Neuro: anesthetics (S1), analgesics (S2), and sedatives (S3), and two for Renal: diuretics and dialysis. Proposed hierarchy is generic which allows integration of agents operating on both continuous and discrete action spaces. However, we tested the approach using discrete action spaces. Dosages for each treatment were discretized into five levels (no-action + 4 quantiles), except dialysis, which is a binary action (active or inactive). Proposed hierarchy supports mixing of any treatment within or across systems, aligning with common clinical practices derived from our data analysis and clinical consultations. For example, we enforced exclusive use of one renal treatment at a time—either diuretics or dialysis, and allowed all other treatment combinations. This showcases the hierarchy’s adaptability to incorporate such domain specific constrains. Accordingly, our solution flexibly handles organ systems with different degrees of complexity and configurations of actions, enabling the effectiveness of the solution in real, complex clinical settings.
RL Reward:
We used a hybrid reward system, combining a mortality-based terminal reward—positive ($+R$) for survival, and negative ($-R$) for death—with clinically guided intermittent rewards to adjust rewards based on immediate health outcomes. This approach addresses reward sparsity in long sequences (see Appendix, Section RL Reward).
Q Learning
Traditional RL frameworks modeled by Markov Decision Processes (MDPs), often assume that actions occur instantaneously with uniform execution times. This assumption simplifies modeling, but fails to capture the complexities of real-world scenarios with actions that span multiple steps and vary in duration.
Therefore we utilize an integration of an options framework with semi-MDPs and decentralized MDPs within a structured hierarchical system. This approach enables an agent to invoke options that ranges from primitive (one-step) to extended (multiple-step) actions at any point. Then the agent regains control to initiate another option only after the previous one completes. Accordingly, the traditional action-value function is adapted to an option-value function $Q^{\pi}(s,o)$ which indicates the value of invoking option $o$ in state $s$ under a policy $\pi$. This is defined as the expected sum of discounted future rewards:
$Q^{\pi}(s,o)=\mathbb{E}\{r_{t+1}+\gamma r_{t+2}+\gamma^{2}r_{t+3}+\dots\mid\epsilon^{\pi}(o,s,t)\}$
where $\gamma$ is a discount factor and $\epsilon^{\pi}(o,s,t)$ is the execution of option $o$ in state $s$ from time $t$ to its termination. Unlike standard options framework, our proposed solution contains both independent and cooperative agents controlling options, as detailed in Hierarchical Decomposition section. Thus, the root level Q-values are updated as follows:
$Q(s,o)\leftarrow Q(s,o)+\alpha\left[Q_{tgt}(s,o)-Q(s,o)\right]$ (1)
where $Q_{tgt}(s,o)$ is determined based on the controlling agent and the option $o$ as follows:
$Q_{tgt}(s,o)=$$\begin{cases}r,&\text{if }o=A_{Rt}^{0}\\ \max\limits_{a\in A_{Neu}}Q_{Neu}(s,a),&\text{if }o=A_{Rt}^{Neu}\\ \max\limits_{a\in A_{Car}}Q_{Car}(s,a),&\text{if }o=A_{Rt}^{Car}\\ \max\limits_{a\in A_{Ren}}Q_{Ren}(s,a),&\text{if }o=A_{Rt}^{Ren}\\ \max\limits_{a\in A_{OMix}}Q_{OMix}(s,a),&\text{if }o=A_{Rt}^{OMix}\end{cases}$
where $r$ is the immediate reward received for choosing no-action ($A_{Rt}^{0}$), and $A_{Neu}$, $A_{Car}$, and $A_{Ren}$ are the action spaces for the Neuro, Cardio and Renal agents, whereas $A_{Rt}^{OMix}$ is their combined action space ($A_{Neu}\times A_{Car}\times A_{Ren}$). Value functions $Q_{Neu}(s,a)$, $Q_{Car}(s,a)$, and $Q_{Ren}(s,a)$ for the corresponding agents are updated via Temporal Difference (TD) learning as follows:
$Q_{Neu}(s,a)\leftarrow$ $Q_{Neu}(s,a)+\alpha\left[y^{Neu}-Q_{Neu}(s,a)\right]$ (2)
$Q_{Car}(s,a)\leftarrow$ $Q_{Car}(s,a)+\alpha\left[y^{Car}-Q_{Car}(s,a)\right]$
$Q_{Ren}(s,a)\leftarrow$ $Q_{Ren}(s,a)+\alpha\left[y^{Ren}-Q_{Ren}(s,a)\right]$
where $y^{X}=r+\gamma\max_{a^{\prime}}Q_{X}(s^{\prime},a^{\prime};\theta^{-})$ for $X\in\{Neu,Car,Ren\}$ are TD targets computed from target networks with parameters $\theta^{-}$. The Q values for each system are updated using observed rewards and estimated values of the next state ($s^{\prime}$).
$M_{OMix}$ is trained using the QMix architecture (Rashid et al. 2020), by combining the individual value functions of its sub-agents ($M_{Omix}^{Neu}$,$M_{Omix}^{Car}$, and $M_{Omix}^{Ren}$) into a unified value function ($Q_{OMix}$). QMix framework supports cooperative training and consistent decision-making across agents by enforcing a monotonicity constraint. The constraint keeps the weights of the mixing network non-negative, ensuring that the $argmax$ operation on $Q_{OMix}$ is consistent with those from the sub-agents. Hence, $Q_{OMix}$ is represented as follows:
$\mathrm{arg\,max}_{a\in A_{OMix}}Q_{OMix}(s,a)=$ (3)
[…]
where $s_{OMix}^{Neu}=\left[s_{t}^{OMix},a_{Car},a_{Ren}\right]$ concatenates the targeted state representation and communicated dosage outputs ($a_{Car}$ and $a_{Ren}$) from the opposite organ-specific agents ($M_{Car}$ and $M_{Ren}$) as described under Hierarchical Decomposition. $s_{OMix}^{Car}$ and $s_{OMix}^{Ren}$ are formed analogously.
Training process:
We used a two phase approach to train the proposed hierarchical model. In phase one, the root agent ($M_{Rt}$) is trained first. Resulting latent embeddings $E^{Rt}$ are then fine-tuned during the subsequent training of organ-specific agents ($M_{Neu}$, $M_{Car}$, and $M_{Ren}$). In the process, targeted embeddings ($E^{Neu}$, $E^{Car}$ and $E^{Ren}$) are learned and formulate Targeted State Representations. These representations are used to train the corresponding lower-level sub-agents. Independent agents directly learn $Q$ values from received rewards (see Equation 2),whereas cooperative agents $M_{Omix}^{Neu}$,$M_{Omix}^{Car}$, $M_{Omix}^{Ren}$ are trained using shared rewards (Equation 3). In phase two, we integrate the full hierarchy by using the trained organ-specific agents from phase one. $Q(s,a)$ values from these agents are used to retrain $M_{Rt}$ (see Equation 1), and no other agent is retrained. After training, at each timestep, $M_{Rt}$ evaluates the patient state $s_{t}^{Rt}$ and select either no-action, or invokes lower level agents for final treatment and dosage recommendations.
Experiments
Dataset
We collected data on 30,440 patients under Sepsis-3 criteria from the popular MIMIC-IV database (Johnson et al. 2020). Data spans from 24 hours pre-diagnosis to 48 hours post-diagnosis forming a maximum 72 hour window per patient, barring death or ICU discharge. Patient state data collected includes 48 physiological measurements including vital signs, laboratory test results, severity scores and demographics (see Appendix, Section Dataset). State and actions data were aggregated into four-hourly windows to generate uniform patient data sequences. The patient trajectories were randomly split into 75% training and 25% testing sets.
Baselines
We evaluate our model performance against a variety of baselines, including both single- (D3QN-S and SoftAC-S) and multi-agent systems under independent (D3QN-O) and cooperative (QMix-O and QMix-T) learning approaches:
  • Clinician: Policy derived from clinician’s recorded action trajectories in the test set.
  • Single D3QN (D3QN-S) (Raghu et al. 2017): State-of-the-art single-agent Dueling DQN predicting all action combinations in a flattened action space. We used the original implementation and tuned hyperparameters, and simplified the problem to only 3 quantiles per treatment.
  • Single SoftAC (SoftAC-S) (Haarnoja et al. 2018): A single-agent Soft Actor-Critic predicting all action combinations in a flattened action space. We used the original implementation and tuned hyperparameters, and simplified the problem to only 3 quantiles per treatment.
  • Organ-specific D3QN (D3QN-O): Three independent D3QN agents for Neuro, Cardio and Renal systems. Models are trained using all available samples, including those treated exclusively for the target organ and those mixed with other organ treatments. We present average quantitative metric values across agents for comparisons.
  • Treatment-specific D3QN (D3QN-T): Six independent D3QN agents for S1, S2, S3, IV, vaso, and diuretics/dialysis. Models are trained using all available samples, including those using single-treatments and mixed-treatments. Average quantitative metrics are reported.
  • Organ-coordinated QMix (QMix-O): Trained end-to-end with three cooperative agents, each corresponding to an organ system operating exclusively within its factored action spaces. Predicts treatments across organs by cooperation, using a QMix mixing network.
  • Treatment-coordinated QMix (QMix-T): Uses six cooperative treatment-level agents (S1, S2, S3, IV fluids, vasopressors, diuretics/dialysis), learning cooperatively via a QMix network.
All models used same state representations, action discretization methods, and reward functions. Detailed feature processing information and model parameters are available in Appendix, Section Model Parameters. All codes for data processing and model training are included in Supplementary materials and will publicly available upon acceptance.
Results and Discussion
We present and discuss our experimental results in the form of answers to four key research questions:
RQ1: Does the proposed solution effectively learns a superior treatment policy?
We evaluate the performance of our learned policies using various off-policy quantitative and qualitative metrics (see Table 1) as follows:
Table 1: Comparison of off-policy evaluation metrics
Model ESS V Mortality (%)
Clinician 89115 18.98 16.27
D3QN-S 12 -1.21 18.69 ± 0.45
SoftAC-S 975 -0.19 18.52 ± 0.57
D3QN-O 99 0.37 14.47 ± 0.34
D3QN-T 373 19.42 14.02 ± 0.41
QMix-O 41 13.80 13.84 ± 0.15
QMix-T 602 25.83 11.29 ± 0.28
Proposed 7233 30.04 8.81 ± 0.24
1) Average Returns: The performance of learned policies is quantitatively evaluated using their estimated average returns. We use $V^{CWPDIS}$ ($V$), the average return from common sub-trajectories in both learned and clinician policies (Thomas 2015). Effective sample size ($ESS$) is the number of common sub-trajectories, capped at the total number in the clinician-policy. $ESS$ measures the confidence in the corresponding $V$ value. An effective policy should have a higher $V$ with a significant $ESS$. Evaluations showcase that all single agent baselines (D3QN-S and SoftAC-S) fail, resulting in negative average returns. This indicates the single agent systems’ inability to learn effective policies in this complex setting. The best performing baseline, QMix-T, had the highest ESS among the baselines and a larger $V$ than the clinician. Thus, while QMix-T’s policy does not fully align with the clinician’s, it can effectively select actions that result in larger immediate returns. Our proposed model considerably outperformed all baselines and the clinician policy, obtaining the highest $V$, supported by a significant $ESS$. This suggests that the proposed model learnt a superior policy which acts in accordance to high-return clinician actions while deviating from low return ones.
2) Mortality Rate: In line with literature, we estimate mortality rates using the clinician policy’s relationship between mortality and expected returns. Specifically, we categorize expected returns from patient trajectories into bins and calculate the mortality rate for each bin. The resulting relationship between the expected returns and mortality is used to estimate mortality rates for learned policies based on their expected returns. The single-agent baselines (D3QN-S and SoftAC-S) both showed an approximate 14% increase in mean mortality compared to the Clinician, indicating a failure to learn an improved policy. In contrast, cooperative multi-agent baselines (QMix-O and QMix-T) demonstrated a decrease of 14.9% and 30.6% in mortality. The non-cooperative multi-agent baselines (D3QN-O and D3QN-T) outperformed single agent baselines, but were less effective than the cooperative multi-agent baselines, highlighting the importance of collaboration among agents to obtain superior policies. Proposed model showcased the highest decrease in mean mortality by 45.9%, highlighting its ability to learn a policy that could considerably improve patient survival compared to both state-of-the-art models and clinician policy.
(a) D3QN-S
(b) SoftAC-S
(c) QMix-O
(d) QMix-T
(e) Proposed
Figure 2: Mortality vs. expected return for all models. The shaded area represents standard errors.
3) Mortality vs. Expected Return: We further evaluate the efficacy of the learned policies by analyzing the correlation between mortality and expected returns (see Figure 2). An effective policy should display a strong negative correlation, where higher expected returns translate to lower mortality, and vice versa. This indicates that the policy learns to associate actions leading to lower returns with higher mortality, and vice versa. The multi-agent baselines (QMix-O and QMix-T) display steeper negative curves than the single-agent baselines (D3QN-S and SoftAC-S). However, these baselines either failed to maintain the negative correlation consistently throughout the plot with some positive correlations in some parts, or did not exhibit a steeper negative correlation. In contrast, the proposed model showed the steepest negative correlation with a consistent negative relationship, indicating the highest ability to enhance patient survival.
Figure 3: Dosage differences (x-axis) versus mortality (y-axis) for the top single-agent baseline SoftAC (top), multi-agent baseline QMix-T (middle), and our proposed model (bottom), with colors blue, red, and green denoting different organ systems.
4) Mortality vs. Difference in Recommended Dosage: For each intervention, we analyzed the relationship between mortality rates, and the recommended dosage differences between clinician-administered and learned policies. We estimated this relationship by categorizing quantile-level dosage differences into bins and computing mortality rates of each bin. An effective policy should align with clinician dosages that resulted in low mortality (x-axis=0), and increasingly deviate from those associate with increasing mortality, ideally forming a V-shaped curve centered at 0.
We showcase the performance of our model against the best-performing single-agent (SoftAC-S) and multi-agent (QMix-T) models (see Figure 3). See Appendix, Section Experimental Results for comparison across all baseline models. The single-agent model failed to achieve the desired V-shape for any treatment, likely due to the large and complex action space. The multi-agent baseline shows comparatively better performance, demonstrating the advantage of cooperative agents operating under factored action spaces. Our model closely approximated the desired V-shape across treatments, except for S3 due to minimal sample sizes (see Appendix, Section RL Action for sample sizes).
Figure 4: Correlation matrices of 7 treatment types within clinician and proposed policies across low, medium, and high SOFA severity levels, categorized by (a) survived and (b) deceased outcomes.
RQ2: Do individual agents learn effective local policies?
In addition to evaluating the overall effectiveness of the learned hierarchical policy, we evaluated the local policies of individual agents within the proposed hierarchy, each operating within their factored state and action spaces. We analyzed mortality rates against clinician and agent recommended dosage differences (see Appendix, Section Experimental Results). All individual agents’ policies closely followed the desired ’V’ shape, indicating effective local policies across the hierarchy. This increases our confidence in the combined global policy and enhances the reliability of our proposed hierarchical task decomposition solution.
RQ3: How does the learned policy compare to the clinician policy in the context of multi-organ treatment decisions?
We assess the alignment between clinician and learned policy using correlations among all treatments within each policy. For granular comparisons, these correlations were further categorized by low, medium and high SOFA severity levels at the time of intervention (see Figure 4). While exact concordance is not expected since survival trajectories could often include non-optimal actions, proposed policy generally aligns with clinician decisions linked to patient survival and deviates from those associated with mortality.
RQ4: How effective are the proposed dual-layer state representations and cross-agent communication mechanisms?
We measured the contributions of two core components of our proposed solution: dual-layer state representations and cross-agent communications. We trained three variations of our hierarchical model: one without inter-agent communications (Prop-NoC), one without the proposed state representations, using raw feature vectors (Prop-NoSR), and without both (Prop-NoC-NoSR). All models showed degraded performance in mortality rates and CWPDIS values (see Tables 1 and 2). Compared to our full proposed model, we see an increase of 19.6%, 27.2%, 33.4% in mean mortality for Prop-NoC, Prop-NoSR, and Prop-NoC-NoSR, respectively. This indicates the critical role of these components in handling the task complexity.
Table 2: Off-policy evaluation metrics of the proposed model, without inter-agent communication and state representations.
Model ESS V Mortality (%)
Prop-NoC 4775 26.27 10.54 ± 0.27
Prop-NoSR 1020 26.13 11.21 ± 0.33
Prop-NoC-NoSR 3633 24.80 11.75 ± 0.50
Conclusions
We introduce a hierarchical multi-agent reinforcement learning framework for the complex and first-of-its-kind multi-organ treatment recommendations, setting a new benchmark in clinical decision support systems. Mimicking real world collaborative clinical settings, our solution effectively decomposes the complex treatment process into a clinically meaningful hierarchy of subtasks. Each subtask is managed by specialized agents operating within dedicated subspaces. It supports both independent and cooperative agent functionality through robust inter-and intra- organ communications. Moreover, a dual-layer state representation technique is proposed to support advanced contextualization needed at multiple levels in the hierarchy. We evaluate our solution on the non-standardized and multi-dimensional sepsis treatment recommendation. Comprehensive quantitative and qualitative evaluation showed that our solution consistently outperformed baselines, significantly improving the patient survival and effectively managing task complexity. Furthermore, learned policy closely followed successful clinical treatment patterns, deviating only when beneficial, thus enhancing the reliability of the policy. The inherent flexibility and scalability of our solution allow it to be expanded to a broader range of treatments and organ systems, and even for complex decision-making scenarios beyond healthcare.
Technical Appendix
Methodology
RL Actions:
For each 4-hour window, we computed the total dosage for each treatment by multiplying its infusion rate with the overlapping duration, and adding any IV push volumes. Treatments with multiple drugs were converted into their standard equivalents (i.e., vasopressors into norepinepherine (Kotani et al. 2023), S2 to fentanyl (McPherson et al. 2010), and diuretics to furosemide (Konerman et al. 2022)) (see Table 1).
Figure 5: Baseline Performances omitted due to space limitations from main text. Dosage differences (x-axis) versus mortality (y-axis) for individual agents in the hierarchy, with colors blue, red, and green denoting different organ systems.
Table 1: Summary of treatments, clinical components, purposes, and number of samples per treatment
Organ System Treatment Clinical Components Purpose Samples
Neurological S1 Propofol Anesthesia 71,389
S2 Fentanyl, Morphine Analgesia (pain relief) 89,400
S3 Dexmedetomidine Sedation 10,789
Cardiovascular IV Fluids Crystalloids, colloids, blood products (tonicity adjusted) Hemodynamic Support, hydration, electrolyte balance 269,672
Vasopressors Norepinephrine, Epinephrine, Dopamine, Vasopressin, Phenylephrine Increase blood pressure for improved organ perfusion; stabilize hemodynamics 55,591
Renal Diuretics Furosemide, Bumetanide Reduce blood pressure and edema; maintain fluid balance 17,279
Dialysis Active renal replacement therapy Waste product removal, maintain electrolyte and fluid balance; support kidney function 7,135
RL Reward:
In addition to the mortality based terminal reward ($+R/-R$), in line with the literature, we use a clinically guided intermittent reward function $R_{im}(s_{t},s_{t+1})$, calculated from the transition from $s_{t}$ to $s_{t+1}$ based on immediate impact on person’s health post-treatment (Raghu et al. 2017).
Table 2: List of model features
Category Feature Name
Demographics (9) Age, Elix., Shock index, SOFA, GCS, Weight, SIRS, Gender, Readmission
Vital signs (10) HR, SBP, MBP, DBP, Resp, Temp., PaCO2, PaO2, PaO2/FiO2 ratio, SpO2
Lab values (24) Albumin, pH, Calcium, Glucose, Hb, Magnesium, WBC, Creatinine, Bicarbonate, Sodium, CO2, Lactate, Chloride, Platelets, Potassium, PTT, PT, AST, ALT, BUN, INR, Ionised calcium, Total bilirubin, Base excess
Output events (2) Fluid output (4 hourly), Total output
Ventilation & others (3) Mechanical ventilation, FiO2, Timestep
  • Abbreviations- INR: International Normalized Ratio; PT: Prothrombin Time; PTT: Partial Thromboplastin Time; SIRS: Systemic Inflammatory Response Syndrome; ICU: Intensive care unit; WBC: White blood cell; Temp.: Temperature; GCS: Glasgow Coma Scale; Resp.: Respiratory rate; HR: Heart rate; SBP: Systolic blood pressure; MBP: Mean blood pressure; DBP: Diastolic blood pressure; Hb: Hemoglobin; Elix.: Elixhauser score.
Experiments
Dataset:
We used 48 physiological features including demographics, lab values, vital signs and intake/output events (see Table 2) and aggregated data into 4-hourly windows using mean or sum as appropriate. Missing values were imputed with the last available data from previous windows. Binary features were normalized to -0.5 and 0.5, and normally and log-normally distributed features to 0-1 range. The study included 25,492 survivors (41.7% female) and 4,948 deceased patients (44.0% female).
Reproducibility and Model Parameters:
Source codes along with tuned parameters and architectures are uploaded as technical appendix with the submission and available online as an anonymous repository [^3] https://anonymous.4open.science/r/HMARL-PyTorch-022B/. All models are tested on the same holdout set. We experimentally set $\gamma$ to 0.99 and $k$ to 8. The sepsis cohort can be replicated using the provided SQL scripts and Python codes.
Mortality vs. Difference in Recommended Dosage:
see Figure 5 for remaining plots omitted from main text.
Individual Agent Performance:
Figure 6 shows mortality verses dosage differences plots of individual agents.
Figure 6: Performances of individual agents in the hierarchy. Dosage differences (x-axis) versus mortality (y-axis) for individual agents in the hierarchy, with colors blue, red, and green denoting different organ systems.
Ethical Considerations:
Similar to RL based Clinical Decision Support Systems (CDSSs) proposed in the literature for various diseases including sepsis management (Komorowski et al. 2018; Raghu et al. 2017; Saria 2018), our solution is a human in the loop CDSS; where clinicians could utilize the data driven decisions provided by the intelligent agents as auxiliary inputs before making the final recommendations. In addition to the final action recommendations from the CDSS, the $Q(s,a)$ values available for all possible actions ($a$) could be used to quantify and compare the quality of the available treatment options for a given patient, with respect to his long term survival. The proposed solution allows clinicians to collaborate with the intelligent agent that combines the experiences of successful clinical decisions in the past, but leaves the full control of the final decision to the clinician. Furthermore, in line with the literature, all models were trained and evaluated using publicly available retrospective data, and no clinical trials were conducted. Therefore, the proposed solution does not raise ethical concerns.